ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 46 Żelazko, Wiesław, Mathematical Institute, Polish Academy of Sciences A supersimple algebra A supersimple algebra is a (real or complex) topological algebra A, dim A > 1, which has no proper closed subalgebras. We shall describe an example of a supersimple algebra (with separately continuous multiplication) given in [1] and discuss some related open questions. Reference: [1] W. Żelazko, A semitopological algebra without proper closed subalgebras, Coll. Math. 74 (1997), 239-242.